OpenStudy (anonymous):

2x/(1-x)^2+2y/(1-y)^2+2z/(1-z)^2=8xyz/((1-x)^2)((1-y)^2)((1-z)^2) when x+y+z=xyz

OpenStudy (studygurl14):

Can you type this up again using the equation thing?

OpenStudy (anonymous):

\[2x/(1-x)^2+2y/(1-y)^2+2z/(1-z)^2=8xyz/((1-x)^2)((1-y)^2)((1-z)^2\]

undeadknight26 (undeadknight26):

dats one long equation...

OpenStudy (studygurl14):

\[\frac{ 2x }{ (1- x)^{2} } + \frac{ 2y }{ (1-y)^{2} } + \frac{ 2x }{ (1-z)^{2} } = \frac{ 8xyz }{ (1 - x)^{2}(1 - y)^{2}(1 - z)^{2} }\] \[x + y + z = xyz\] Hm...this certainly is an interesting equation. What exactly are you trying to find. I'm confused about that.

OpenStudy (anonymous):

you can use tana=x tanb=y tanb=z

OpenStudy (studygurl14):

?